OpenStudy (anonymous):

Show that a line through (0, 0) and (c, d) is perpendicular to a line through (0, 0) and (-d, c). Slope of first line = Slope of second line = REALLY DONT UNDERSTAND THIS...

OpenStudy (saifoo.khan):

Slopes..

OpenStudy (anonymous):

how can i do it with numbers!!!!!!!

OpenStudy (anonymous):

its harder

OpenStudy (saifoo.khan):

for perpendicular lines, use: m1 * m2 = -1

OpenStudy (saifoo.khan):

Like let's say the slope of a line is 2. then the perpendicular line will have, m1 * m2 = -1 2 * m2 = -1 m2 = -1/2

OpenStudy (anonymous):

so.. let me guess the answers are either c/d or -d/c

OpenStudy (precal):

 slope of first line